# Continuous Diffusion for Mixed-Type Tabular Data

### v2

@@ Appendix I Machine Learning Efficiency Models
ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
acsincome $0.808_{\pm 0.001}$ $0.989_{\pm 0.001}$ $0.985_{\pm 0.000}$ $0.825_{\pm 0.002}$ $0.540_{\pm 0.003}$ $0.532_{\pm 0.004}$ $0.526_{\pm 0.002}$
adult $0.889_{\pm 0.002}$ $0.997_{\pm 0.000}$ $0.967_{\pm 0.001}$ $0.590_{\pm 0.003}$ $0.992_{\pm 0.001}$ $0.604_{\pm 0.002}$ $0.588_{\pm 0.002}$ $0.591_{\pm 0.005}$
bank $0.955_{\pm 0.002}$ $1.000_{\pm 0.000}$ $0.988_{\pm 0.001}$ $0.783_{\pm 0.003}$ $1.000_{\pm 0.000}$ $0.795_{\pm 0.003}$ $0.739_{\pm 0.003}$ $0.694_{\pm 0.006}$
beijing $0.995_{\pm 0.000}$ $0.998_{\pm 0.000}$ $0.993_{\pm 0.001}$ $0.966_{\pm 0.002}$ $0.997_{\pm 0.001}$ $0.951_{\pm 0.002}$ $0.949_{\pm 0.001}$ $0.947_{\pm 0.002}$
churn $0.853_{\pm 0.002}$ $0.945_{\pm 0.006}$ $0.843_{\pm 0.011}$ $0.561_{\pm 0.005}$ $0.730_{\pm 0.012}$ $0.621_{\pm 0.016}$ $0.533_{\pm 0.007}$ $0.544_{\pm 0.031}$
covertype $0.945_{\pm 0.002}$ $0.997_{\pm 0.000}$ $0.989_{\pm 0.001}$ $0.586_{\pm 0.002}$ $0.900_{\pm 0.002}$ $0.991_{\pm 0.001}$ $0.992_{\pm 0.001}$ $0.991_{\pm 0.001}$
default $0.991_{\pm 0.001}$ $0.998_{\pm 0.001}$ $0.997_{\pm 0.001}$ $0.821_{\pm 0.002}$ $0.995_{\pm 0.000}$ $0.827_{\pm 0.004}$ $0.802_{\pm 0.003}$ $0.871_{\pm 0.001}$
diabetes $0.854_{\pm 0.002}$ $0.935_{\pm 0.002}$ $0.997_{\pm 0.001}$ $0.858_{\pm 0.001}$ $0.780_{\pm 0.002}$ $0.866_{\pm 0.002}$
lending $0.997_{\pm 0.001}$ $0.995_{\pm 0.002}$ $0.995_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.955_{\pm 0.006}$ $0.954_{\pm 0.009}$ $0.961_{\pm 0.004}$
news $0.998_{\pm 0.000}$ $1.000_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.966_{\pm 0.002}$ $1.000_{\pm 0.000}$ $0.973_{\pm 0.001}$ $0.953_{\pm 0.001}$ $0.977_{\pm 0.001}$
nmes $0.987_{\pm 0.002}$ $0.992_{\pm 0.003}$ $0.988_{\pm 0.002}$ $0.650_{\pm 0.014}$ $0.988_{\pm 0.000}$ $0.636_{\pm 0.008}$ $0.623_{\pm 0.008}$ $0.642_{\pm 0.010}$
@@
Table 11: Distance to closest record of the generated data (incl. standard errors in subscripts) for five benchmark models and for CDTD with three different noise schedules.
@@
ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
acsincome $8.637_{\pm 0.027}$ $10.758_{\pm 0.054}$ $6.652_{\pm 0.032}$ $10.877_{\pm 0.092}$ $8.346_{\pm 0.056}$ $8.322_{\pm 0.047}$ $8.349_{\pm 0.033}$
adult $2.523_{\pm 0.012}$ $5.012_{\pm 0.028}$ $2.227_{\pm 0.013}$ $1.647_{\pm 0.009}$ $2.735_{\pm 0.028}$ $1.112_{\pm 0.019}$ $1.231_{\pm 0.011}$ $1.294_{\pm 0.009}$
bank $3.025_{\pm 0.017}$ $3.840_{\pm 0.014}$ $3.136_{\pm 0.007}$ $2.327_{\pm 0.010}$ $3.062_{\pm 0.012}$ $1.828_{\pm 0.008}$ $1.943_{\pm 0.007}$ $2.062_{\pm 0.008}$
beijing $0.735_{\pm 0.003}$ $1.004_{\pm 0.006}$ $0.926_{\pm 0.003}$ $0.739_{\pm 0.006}$ $0.610_{\pm 0.002}$ $0.490_{\pm 0.002}$ $0.489_{\pm 0.001}$ $0.477_{\pm 0.002}$
churn $1.136_{\pm 0.015}$ $1.804_{\pm 0.036}$ $1.146_{\pm 0.039}$ $0.342_{\pm 0.031}$ $0.852_{\pm 0.016}$ $0.332_{\pm 0.021}$ $0.274_{\pm 0.021}$ $0.276_{\pm 0.012}$
covertype $1.741_{\pm 0.011}$ $5.773_{\pm 0.017}$ $3.173_{\pm 0.013}$ $0.889_{\pm 0.007}$ $1.508_{\pm 0.020}$ $2.297_{\pm 0.026}$ $2.209_{\pm 0.022}$ $2.252_{\pm 0.013}$
default $3.095_{\pm 0.026}$ $5.880_{\pm 0.020}$ $3.216_{\pm 0.013}$ $1.422_{\pm 0.013}$ $2.593_{\pm 0.020}$ $1.127_{\pm 0.028}$ $1.269_{\pm 0.014}$ $1.253_{\pm 0.012}$
diabetes $17.736_{\pm 0.107}$ $21.935_{\pm 0.046}$ $8.214_{\pm 0.022}$ $15.279_{\pm 0.026}$ $15.126_{\pm 0.058}$ $15.350_{\pm 0.059}$
lending $17.776_{\pm 0.132}$ $20.239_{\pm 0.222}$ $10.688_{\pm 0.025}$ $12.537_{\pm 0.076}$ $13.775_{\pm 0.147}$ $14.162_{\pm 0.188}$ $13.966_{\pm 0.282}$
news $6.147_{\pm 0.010}$ $4.789_{\pm 0.005}$ $5.821_{\pm 0.003}$ $4.960_{\pm 0.006}$ $4.661_{\pm 0.023}$ $3.635_{\pm 0.004}$ $3.687_{\pm 0.006}$ $3.749_{\pm 0.048}$
nmes $2.203_{\pm 0.028}$ $2.971_{\pm 0.008}$ $1.710_{\pm 0.019}$ $0.891_{\pm 0.033}$ $1.231_{\pm 0.024}$ $0.664_{\pm 0.029}$ $0.710_{\pm 0.032}$ $0.771_{\pm 0.023}$
@@
Table 12: Machine learning efficiency F1 score for five benchmark models, the real training data and for CDTD with three different noise schedules. The standard deviation takes into account five different sampling seeds and uses the average results of the four machine learning efficiency models computed across ten model seeds.
@@
Real Data ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
adult $0.797_{\pm 0.000}$ $0.769_{\pm 0.002}$ $0.647_{\pm 0.015}$ $0.756_{\pm 0.002}$ $0.787_{\pm 0.001}$ $0.745_{\pm 0.004}$ $0.787_{\pm 0.001}$ $0.787_{\pm 0.001}$ $0.787_{\pm 0.001}$
bank $0.745_{\pm 0.002}$ $0.682_{\pm 0.006}$ $0.680_{\pm 0.006}$ $0.629_{\pm 0.006}$ $0.720_{\pm 0.006}$ $0.673_{\pm 0.006}$ $0.776_{\pm 0.003}$ $0.767_{\pm 0.004}$ $0.737_{\pm 0.004}$
churn $0.873_{\pm 0.003}$ $0.780_{\pm 0.015}$ $0.761_{\pm 0.009}$ $0.802_{\pm 0.017}$ $0.857_{\pm 0.007}$ $0.865_{\pm 0.008}$ $0.854_{\pm 0.011}$ $0.852_{\pm 0.006}$ $0.845_{\pm 0.011}$
covertype $0.817_{\pm 0.001}$ $0.783_{\pm 0.001}$ $0.442_{\pm 0.008}$ $0.711_{\pm 0.002}$ $0.799_{\pm 0.001}$ $0.767_{\pm 0.001}$ $0.734_{\pm 0.002}$ $0.754_{\pm 0.001}$ $0.722_{\pm 0.002}$
default $0.674_{\pm 0.001}$ $0.627_{\pm 0.003}$ $0.686_{\pm 0.002}$ $0.632_{\pm 0.007}$ $0.678_{\pm 0.002}$ $0.638_{\pm 0.008}$ $0.670_{\pm 0.002}$ $0.671_{\pm 0.001}$ $0.673_{\pm 0.003}$
diabetes $0.621_{\pm 0.002}$ $0.572_{\pm 0.005}$ $0.557_{\pm 0.004}$ $0.553_{\pm 0.003}$ $0.617_{\pm 0.002}$ $0.617_{\pm 0.002}$ $0.611_{\pm 0.002}$
@@
Table 13: Machine learning efficiency AUC score for five benchmark models, the real training data and for CDTD with three different noise schedules. The standard deviation takes into account five different sampling seeds and uses the average results of the four machine learning efficiency models computed across ten model seeds.
@@
Real Data ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
adult $0.915_{\pm 0.000}$ $0.901_{\pm 0.000}$ $0.836_{\pm 0.006}$ $0.889_{\pm 0.002}$ $0.908_{\pm 0.000}$ $0.880_{\pm 0.005}$ $0.910_{\pm 0.000}$ $0.910_{\pm 0.001}$ $0.909_{\pm 0.000}$
bank $0.947_{\pm 0.000}$ $0.938_{\pm 0.001}$ $0.934_{\pm 0.003}$ $0.830_{\pm 0.020}$ $0.940_{\pm 0.005}$ $0.929_{\pm 0.005}$ $0.945_{\pm 0.000}$ $0.945_{\pm 0.001}$ $0.943_{\pm 0.004}$
churn $0.964_{\pm 0.001}$ $0.939_{\pm 0.007}$ $0.882_{\pm 0.006}$ $0.948_{\pm 0.004}$ $0.957_{\pm 0.004}$ $0.961_{\pm 0.001}$ $0.962_{\pm 0.001}$ $0.962_{\pm 0.001}$ $0.959_{\pm 0.003}$
covertype $0.892_{\pm 0.000}$ $0.860_{\pm 0.001}$ $0.677_{\pm 0.007}$ $0.777_{\pm 0.001}$ $0.876_{\pm 0.000}$ $0.845_{\pm 0.001}$ $0.816_{\pm 0.002}$ $0.828_{\pm 0.001}$ $0.802_{\pm 0.002}$
default $0.768_{\pm 0.000}$ $0.754_{\pm 0.002}$ $0.744_{\pm 0.002}$ $0.751_{\pm 0.004}$ $0.763_{\pm 0.002}$ $0.739_{\pm 0.008}$ $0.762_{\pm 0.003}$ $0.765_{\pm 0.002}$ $0.765_{\pm 0.002}$
diabetes $0.693_{\pm 0.001}$ $0.669_{\pm 0.002}$ $0.626_{\pm 0.003}$ $0.592_{\pm 0.002}$ $0.675_{\pm 0.001}$ $0.673_{\pm 0.001}$ $0.667_{\pm 0.001}$
@@
Table 14: Machine learning efficiency RMSE for five benchmark models, the real training data and for CDTD with three different noise schedules. The standard deviation takes into account five different sampling seeds and uses the average results of the four machine learning efficiency models computed across ten model seeds.
@@
Real Data ARF CTGAN TVAE TabDDPM CoDi CDTD (single) CDTD (per type) CDTD (per feature)
acsincome $0.804_{\pm 0.012}$ $0.757_{\pm 0.007}$ $2.292_{\pm 0.013}$ $1.054_{\pm 0.011}$ $0.857_{\pm 0.010}$ $0.838_{\pm 0.015}$ $0.811_{\pm 0.014}$ $0.820_{\pm 0.011}$
beijing $0.712_{\pm 0.001}$ $0.792_{\pm 0.007}$ $1.246_{\pm 0.010}$ $1.690_{\pm 0.016}$ $0.606_{\pm 0.006}$ $0.912_{\pm 0.005}$ $0.774_{\pm 0.005}$ $0.770_{\pm 0.005}$ $0.762_{\pm 0.005}$
lending $0.030_{\pm 0.000}$ $0.274_{\pm 0.007}$ $0.137_{\pm 0.007}$ $0.404_{\pm 0.007}$ $0.795_{\pm 0.031}$ $0.061_{\pm 0.001}$ $0.060_{\pm 0.001}$ $0.066_{\pm 0.002}$
news $1.001_{\pm 0.002}$ $0.923_{\pm 0.052}$ $1.906_{\pm 0.019}$ $3.999_{\pm 0.175}$ $0.083_{\pm 0.001}$ $1.302_{\pm 0.074}$ $0.819_{\pm 0.103}$ $0.776_{\pm 0.091}$ $0.755_{\pm 0.066}$
nmes $1.001_{\pm 0.003}$ $0.972_{\pm 0.024}$ $1.331_{\pm 0.052}$ $1.127_{\pm 0.047}$ $1.154_{\pm 0.047}$ $1.137_{\pm 0.052}$ $1.108_{\pm 0.083}$ $1.184_{\pm 0.076}$ $1.203_{\pm 0.081}$
Appendix O Ablation Study Details
@@
Appendix P Training and Sampling Times Details
@@
ARF CTGAN TVAE TabDDPM CoDi CDTD (per feature)
acsincome 80.3 59.9 26.0 - 231.9 5.8
adult 7.4 36.2 23.7 38.3 48.3 6.9
bank 11.0 37.6 24.6 40.5 42.7 26.3
beijing 3.7 34.3 23.9 36.1 24.9 23.4
churn 0.3 27.1 13.7 18.2 25.7 6.1
covertype 130.2 58.0 36.5 44.9 69.2 28.2
default 12.0 38.3 24.8 38.9 45.9 26.4
diabetes 58.5 90.1 25.3 - 870 (est.) 26.9
lending 5.2 157.9 36.6 48.7 3000 (est.) 25.3
news 23.0 48.8 33.3 37.2 41.5 25.2
nmes 0.4 32.8 17.2 24.9 30.2 6.3
@@
ARF CTGAN TVAE TabDDPM CoDi CDTD (per feature)
acsincome 4.20 0.23 0.07 - 10.26 0.59
adult 1.78 0.31 0.16 0.82 3.65 0.56
bank 2.24 0.44 0.44 0.87 3.38 0.64
beijing 0.34 0.41 0.32 2.09 2.45 0.26
churn 1.00 0.40 0.24 0.95 2.78 0.39
covertype 9.74 0.28 0.25 2.45 4.35 1.97
default 2.07 0.27 0.25 0.86 3.48 0.60
diabetes 5.87 0.53 0.15 - - 1.33
lending 2.49 0.45 0.54 4.33 - 0.69
news 3.89 0.43 0.30 5.13 2.93 0.85
nmes 1.54 0.31 0.17 4.17 2.91 0.55

### v3

@@ Appendix I Machine Learning Efficiency Models
SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
acsincome - $0.808_{\pm 0.001}$ $0.989_{\pm 0.001}$ $0.985_{\pm 0.000}$ $0.825_{\pm 0.002}$ $0.680_{\pm 0.002}$ $0.540_{\pm 0.003}$ $0.532_{\pm 0.004}$ $0.526_{\pm 0.002}$
adult $0.320_{\pm 0.006}$ $0.889_{\pm 0.002}$ $0.997_{\pm 0.000}$ $0.967_{\pm 0.001}$ $0.590_{\pm 0.003}$ $0.992_{\pm 0.001}$ $0.630_{\pm 0.003}$ $0.604_{\pm 0.002}$ $0.588_{\pm 0.002}$ $0.591_{\pm 0.005}$
bank $0.633_{\pm 0.008}$ $0.955_{\pm 0.002}$ $1.000_{\pm 0.000}$ $0.988_{\pm 0.001}$ $0.783_{\pm 0.003}$ $1.000_{\pm 0.000}$ $0.843_{\pm 0.002}$ $0.795_{\pm 0.003}$ $0.739_{\pm 0.003}$ $0.694_{\pm 0.006}$
beijing $0.976_{\pm 0.001}$ $0.995_{\pm 0.000}$ $0.998_{\pm 0.000}$ $0.993_{\pm 0.001}$ $0.966_{\pm 0.002}$ $0.997_{\pm 0.001}$ $0.966_{\pm 0.001}$ $0.951_{\pm 0.002}$ $0.949_{\pm 0.001}$ $0.947_{\pm 0.002}$
churn $0.339_{\pm 0.020}$ $0.853_{\pm 0.002}$ $0.945_{\pm 0.006}$ $0.843_{\pm 0.011}$ $0.561_{\pm 0.005}$ $0.730_{\pm 0.012}$ $0.865_{\pm 0.012}$ $0.621_{\pm 0.016}$ $0.533_{\pm 0.007}$ $0.544_{\pm 0.031}$
covertype - $0.945_{\pm 0.002}$ $0.997_{\pm 0.000}$ $0.989_{\pm 0.001}$ $0.586_{\pm 0.002}$ $0.900_{\pm 0.002}$ $0.979_{\pm 0.001}$ $0.991_{\pm 0.001}$ $0.992_{\pm 0.001}$ $0.991_{\pm 0.001}$
default $0.493_{\pm 0.009}$ $0.991_{\pm 0.001}$ $0.998_{\pm 0.001}$ $0.997_{\pm 0.001}$ $0.821_{\pm 0.002}$ $0.995_{\pm 0.000}$ $0.902_{\pm 0.001}$ $0.827_{\pm 0.004}$ $0.802_{\pm 0.003}$ $0.871_{\pm 0.001}$
diabetes $0.367_{\pm 0.001}$ $0.854_{\pm 0.002}$ $0.935_{\pm 0.002}$ $0.997_{\pm 0.001}$ $0.940_{\pm 0.001}$ $0.858_{\pm 0.001}$ $0.780_{\pm 0.002}$ $0.866_{\pm 0.002}$
lending $0.926_{\pm 0.004}$ $0.997_{\pm 0.001}$ $0.995_{\pm 0.002}$ $0.995_{\pm 0.001}$ $1.000_{\pm 0.000}$ $0.998_{\pm 0.001}$ $0.955_{\pm 0.006}$ $0.954_{\pm 0.009}$ $0.961_{\pm 0.004}$
news $0.993_{\pm 0.001}$ $0.998_{\pm 0.000}$ $1.000_{\pm 0.000}$ $1.000_{\pm 0.000}$ $0.966_{\pm 0.002}$ $1.000_{\pm 0.000}$ $0.999_{\pm 0.000}$ $0.973_{\pm 0.001}$ $0.953_{\pm 0.001}$ $0.977_{\pm 0.001}$
nmes $0.905_{\pm 0.007}$ $0.987_{\pm 0.002}$ $0.992_{\pm 0.003}$ $0.988_{\pm 0.002}$ $0.650_{\pm 0.014}$ $0.988_{\pm 0.000}$ $0.841_{\pm 0.008}$ $0.636_{\pm 0.008}$ $0.623_{\pm 0.008}$ $0.642_{\pm 0.010}$
@@
Table 16: Distance to closest record of the generated data (incl. standard errors in subscripts) for seven benchmark models and for CDTD with three different noise schedules.
@@
Test Set SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
acsincome $7.673_{\pm 0.017}$ $8.637_{\pm 0.027}$ $10.758_{\pm 0.054}$ $6.652_{\pm 0.032}$ $10.877_{\pm 0.092}$ $10.305_{\pm 0.073}$ $8.346_{\pm 0.056}$ $8.322_{\pm 0.047}$ $8.349_{\pm 0.033}$
adult $1.870_{\pm 0.000}$ $1.371_{\pm 0.018}$ $2.523_{\pm 0.012}$ $5.012_{\pm 0.028}$ $2.227_{\pm 0.013}$ $1.647_{\pm 0.009}$ $2.735_{\pm 0.028}$ $2.341_{\pm 0.013}$ $1.112_{\pm 0.019}$ $1.231_{\pm 0.011}$ $1.294_{\pm 0.009}$
bank $2.369_{\pm 0.000}$ $1.369_{\pm 0.011}$ $3.025_{\pm 0.017}$ $3.840_{\pm 0.014}$ $3.136_{\pm 0.007}$ $2.327_{\pm 0.010}$ $3.062_{\pm 0.012}$ $2.973_{\pm 0.012}$ $1.828_{\pm 0.008}$ $1.943_{\pm 0.007}$ $2.062_{\pm 0.008}$
beijing $0.385_{\pm 0.000}$ $0.139_{\pm 0.003}$ $0.735_{\pm 0.003}$ $1.004_{\pm 0.006}$ $0.926_{\pm 0.003}$ $0.739_{\pm 0.006}$ $0.610_{\pm 0.002}$ $0.626_{\pm 0.001}$ $0.490_{\pm 0.002}$ $0.489_{\pm 0.001}$ $0.477_{\pm 0.002}$
churn $0.347_{\pm 0.000}$ $0.232_{\pm 0.028}$ $1.136_{\pm 0.015}$ $1.804_{\pm 0.036}$ $1.146_{\pm 0.039}$ $0.342_{\pm 0.031}$ $0.852_{\pm 0.016}$ $1.130_{\pm 0.018}$ $0.332_{\pm 0.021}$ $0.274_{\pm 0.021}$ $0.276_{\pm 0.012}$
covertype $0.529_{\pm 0.001}$ $1.741_{\pm 0.011}$ $5.773_{\pm 0.017}$ $3.173_{\pm 0.013}$ $0.889_{\pm 0.007}$ $1.508_{\pm 0.020}$ $3.086_{\pm 0.009}$ $2.297_{\pm 0.026}$ $2.209_{\pm 0.022}$ $2.252_{\pm 0.013}$
default $1.812_{\pm 0.000}$ $1.032_{\pm 0.010}$ $3.095_{\pm 0.026}$ $5.880_{\pm 0.020}$ $3.216_{\pm 0.013}$ $1.422_{\pm 0.013}$ $2.593_{\pm 0.020}$ $2.603_{\pm 0.018}$ $1.127_{\pm 0.028}$ $1.269_{\pm 0.014}$ $1.253_{\pm 0.012}$
diabetes $15.608_{\pm 0.055}$ $13.909_{\pm 0.050}$ $17.736_{\pm 0.107}$ $21.935_{\pm 0.046}$ $8.214_{\pm 0.022}$ $28.955_{\pm 0.060}$ $15.279_{\pm 0.026}$ $15.126_{\pm 0.058}$ $15.350_{\pm 0.059}$
lending $11.184_{\pm 0.000}$ $17.752_{\pm 0.143}$ $17.776_{\pm 0.132}$ $20.239_{\pm 0.222}$ $10.688_{\pm 0.025}$ $12.537_{\pm 0.076}$ $16.222_{\pm 0.092}$ $13.775_{\pm 0.147}$ $14.162_{\pm 0.188}$ $13.966_{\pm 0.282}$
news $3.615_{\pm 0.000}$ $3.553_{\pm 0.134}$ $6.147_{\pm 0.010}$ $4.789_{\pm 0.005}$ $5.821_{\pm 0.003}$ $4.960_{\pm 0.006}$ $4.661_{\pm 0.023}$ $5.351_{\pm 0.008}$ $3.635_{\pm 0.004}$ $3.687_{\pm 0.006}$ $3.749_{\pm 0.048}$
nmes $1.931_{\pm 0.000}$ $1.394_{\pm 0.019}$ $2.203_{\pm 0.028}$ $2.971_{\pm 0.008}$ $1.710_{\pm 0.019}$ $0.891_{\pm 0.033}$ $1.231_{\pm 0.024}$ $2.260_{\pm 0.034}$ $0.664_{\pm 0.029}$ $0.710_{\pm 0.032}$ $0.771_{\pm 0.023}$
@@
Table 17: Machine learning efficiency F1 score for seven benchmark models, the real training data and for CDTD with three different noise schedules. The standard deviation takes into account five different sampling seeds and uses the average results of the four machine learning efficiency models computed across ten model seeds.
@@
Real Data SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
adult $0.797_{\pm 0.000}$ $0.784_{\pm 0.001}$ $0.769_{\pm 0.002}$ $0.647_{\pm 0.015}$ $0.756_{\pm 0.002}$ $0.787_{\pm 0.001}$ $0.745_{\pm 0.004}$ $0.782_{\pm 0.001}$ $0.787_{\pm 0.001}$ $0.787_{\pm 0.001}$ $0.787_{\pm 0.001}$
bank $0.745_{\pm 0.002}$ $0.740_{\pm 0.004}$ $0.682_{\pm 0.006}$ $0.680_{\pm 0.006}$ $0.629_{\pm 0.006}$ $0.720_{\pm 0.006}$ $0.673_{\pm 0.006}$ $0.711_{\pm 0.007}$ $0.776_{\pm 0.003}$ $0.767_{\pm 0.004}$ $0.737_{\pm 0.004}$
churn $0.873_{\pm 0.003}$ $0.865_{\pm 0.008}$ $0.780_{\pm 0.015}$ $0.761_{\pm 0.009}$ $0.802_{\pm 0.017}$ $0.857_{\pm 0.007}$ $0.865_{\pm 0.008}$ $0.771_{\pm 0.014}$ $0.854_{\pm 0.011}$ $0.852_{\pm 0.006}$ $0.845_{\pm 0.011}$
covertype $0.817_{\pm 0.001}$ $0.783_{\pm 0.001}$ $0.442_{\pm 0.008}$ $0.711_{\pm 0.002}$ $0.799_{\pm 0.001}$ $0.767_{\pm 0.001}$ $0.614_{\pm 0.015}$ $0.734_{\pm 0.002}$ $0.754_{\pm 0.001}$ $0.722_{\pm 0.002}$
default $0.674_{\pm 0.001}$ $0.677_{\pm 0.001}$ $0.627_{\pm 0.003}$ $0.686_{\pm 0.002}$ $0.632_{\pm 0.007}$ $0.678_{\pm 0.002}$ $0.638_{\pm 0.008}$ $0.496_{\pm 0.009}$ $0.670_{\pm 0.002}$ $0.671_{\pm 0.001}$ $0.673_{\pm 0.003}$
diabetes $0.621_{\pm 0.002}$ $0.615_{\pm 0.002}$ $0.572_{\pm 0.005}$ $0.557_{\pm 0.004}$ $0.553_{\pm 0.003}$ $0.560_{\pm 0.006}$ $0.617_{\pm 0.002}$ $0.617_{\pm 0.002}$ $0.611_{\pm 0.002}$
@@
Table 18: Machine learning efficiency AUC score for seven benchmark models, the real training data and for CDTD with three different noise schedules. The standard deviation takes into account five different sampling seeds and uses the average results of the four machine learning efficiency models computed across ten model seeds.
@@
Real Data SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
adult $0.915_{\pm 0.000}$ $0.906_{\pm 0.001}$ $0.901_{\pm 0.000}$ $0.836_{\pm 0.006}$ $0.889_{\pm 0.002}$ $0.908_{\pm 0.000}$ $0.880_{\pm 0.005}$ $0.906_{\pm 0.001}$ $0.910_{\pm 0.000}$ $0.910_{\pm 0.001}$ $0.909_{\pm 0.000}$
bank $0.947_{\pm 0.000}$ $0.943_{\pm 0.001}$ $0.938_{\pm 0.001}$ $0.934_{\pm 0.003}$ $0.830_{\pm 0.020}$ $0.940_{\pm 0.005}$ $0.929_{\pm 0.005}$ $0.939_{\pm 0.003}$ $0.945_{\pm 0.000}$ $0.945_{\pm 0.001}$ $0.943_{\pm 0.004}$
churn $0.964_{\pm 0.001}$ $0.961_{\pm 0.002}$ $0.939_{\pm 0.007}$ $0.882_{\pm 0.006}$ $0.948_{\pm 0.004}$ $0.957_{\pm 0.004}$ $0.961_{\pm 0.001}$ $0.919_{\pm 0.006}$ $0.962_{\pm 0.001}$ $0.962_{\pm 0.001}$ $0.959_{\pm 0.003}$
covertype $0.892_{\pm 0.000}$ $0.860_{\pm 0.001}$ $0.677_{\pm 0.007}$ $0.777_{\pm 0.001}$ $0.876_{\pm 0.000}$ $0.845_{\pm 0.001}$ $0.671_{\pm 0.013}$ $0.816_{\pm 0.002}$ $0.828_{\pm 0.001}$ $0.802_{\pm 0.002}$
default $0.768_{\pm 0.000}$ $0.759_{\pm 0.003}$ $0.754_{\pm 0.002}$ $0.744_{\pm 0.002}$ $0.751_{\pm 0.004}$ $0.763_{\pm 0.002}$ $0.739_{\pm 0.008}$ $0.746_{\pm 0.011}$ $0.762_{\pm 0.003}$ $0.765_{\pm 0.002}$ $0.765_{\pm 0.002}$
diabetes $0.693_{\pm 0.001}$ $0.679_{\pm 0.001}$ $0.669_{\pm 0.002}$ $0.626_{\pm 0.003}$ $0.592_{\pm 0.002}$ $0.645_{\pm 0.002}$ $0.675_{\pm 0.001}$ $0.673_{\pm 0.001}$ $0.667_{\pm 0.001}$
@@
Table 19: Machine learning efficiency RMSE for seven benchmark models, the real training data and for CDTD with three different noise schedules. The standard deviation takes into account five different sampling seeds and uses the average results of the four machine learning efficiency models computed across ten model seeds.
@@
Real Data SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (single) CDTD (per type) CDTD (per feature)
acsincome $0.804_{\pm 0.012}$ $0.757_{\pm 0.007}$ $2.292_{\pm 0.013}$ $1.054_{\pm 0.011}$ $0.857_{\pm 0.010}$ $0.959_{\pm 0.012}$ $0.838_{\pm 0.015}$ $0.811_{\pm 0.014}$ $0.820_{\pm 0.011}$
beijing $0.712_{\pm 0.001}$ $0.739_{\pm 0.002}$ $0.792_{\pm 0.007}$ $1.246_{\pm 0.010}$ $1.690_{\pm 0.016}$ $0.606_{\pm 0.006}$ $0.912_{\pm 0.005}$ $0.788_{\pm 0.011}$ $0.774_{\pm 0.005}$ $0.770_{\pm 0.005}$ $0.762_{\pm 0.005}$
lending $0.030_{\pm 0.000}$ $0.042_{\pm 0.001}$ $0.274_{\pm 0.007}$ $0.137_{\pm 0.007}$ $0.404_{\pm 0.007}$ $0.795_{\pm 0.031}$ $0.268_{\pm 0.004}$ $0.061_{\pm 0.001}$ $0.060_{\pm 0.001}$ $0.066_{\pm 0.002}$
news $1.001_{\pm 0.002}$ $1.180_{\pm 0.107}$ $0.923_{\pm 0.052}$ $1.906_{\pm 0.019}$ $3.999_{\pm 0.175}$ $0.083_{\pm 0.001}$ $1.302_{\pm 0.074}$ $0.374_{\pm 0.028}$ $0.819_{\pm 0.103}$ $0.776_{\pm 0.091}$ $0.755_{\pm 0.066}$
nmes $1.001_{\pm 0.003}$ $1.112_{\pm 0.044}$ $0.972_{\pm 0.024}$ $1.331_{\pm 0.052}$ $1.127_{\pm 0.047}$ $1.154_{\pm 0.047}$ $1.137_{\pm 0.052}$ $0.535_{\pm 0.013}$ $1.108_{\pm 0.083}$ $1.184_{\pm 0.076}$ $1.203_{\pm 0.081}$
Appendix S Ablation Study Details
@@
Appendix T Training and Sampling Times Details
@@
SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (per feature)
acsincome - 80.3 59.9 26.0 - 231.9 13.4 5.8
adult - 7.4 36.2 23.7 38.3 48.3 32.7 6.9
bank - 11.0 37.6 24.6 40.5 42.7 48.5 26.3
beijing - 3.7 34.3 23.9 36.1 24.9 25.8 23.4
churn - 0.3 27.1 13.7 18.2 25.7 21.5 6.1
covertype - 130.2 58.0 36.5 44.9 69.2 30.7 28.2
default - 12.0 38.3 24.8 38.9 45.9 40.1 26.4
diabetes - 58.5 90.1 25.3 - 870 (est.) 34.6 26.9
lending - 5.2 157.9 36.6 48.7 3000 (est.) 42.1 25.3
news - 23.0 48.8 33.3 37.2 41.5 57.9 25.2
nmes - 0.4 32.8 17.2 24.9 30.2 31.0 6.3
@@
SMOTE ARF CTGAN TVAE TabDDPM CoDi TabSyn CDTD (per feature)
acsincome 4674.45 4.20 0.23 0.07 - 10.26 3.53 0.59
adult 10.71 1.78 0.31 0.16 0.82 3.65 0.88 0.56
bank 16.19 2.24 0.44 0.44 0.87 3.38 0.80 0.64
beijing 3.98 0.34 0.41 0.32 2.09 2.45 0.99 0.26
churn 0.52 1.00 0.40 0.24 0.95 2.78 0.80 0.39
covertype 10913.34 9.74 0.28 0.25 2.45 4.35 0.85 1.97
default 10.00 2.07 0.27 0.25 0.86 3.48 0.82 0.60
diabetes 166.75 5.87 0.53 0.15 - - 0.83 1.33
lending 4.06 2.49 0.45 0.54 4.33 - 0.85 0.69
news 66.49 3.89 0.43 0.30 5.13 2.93 0.86 0.85
nmes 0.69 1.54 0.31 0.17 4.17 2.91 0.82 0.55